OpenStudy (anonymous):

From a boat on the lake, the angle of elevation to the top of a cliff is 11°50'. If the base of the cliff is 1568 feet from the boat, how high is the cliff (to the nearest foot)?

OpenStudy (anonymous):

help help!

OpenStudy (anonymous):

\[Tan(11+\frac{ 5 }{ 6 })=\frac{ x }{ 1568}\] Can you solve from here?

OpenStudy (anonymous):

@evansda ummm im not sure.. can you explain little further?

OpenStudy (mathteacher1729):

Did you draw a picture? Understanding the "angle of elevation" is crucial.

OpenStudy (anonymous):

 then use the following \[\Large \frac{h}{1568}=\tan(11.08)\]